OpenStudy (anonymous):

Please Help Me with a radical problem:

OpenStudy (anonymous):

\[\sqrt 140x^3\]

OpenStudy (anonymous):

Can You Help Me Please???????????????????

OpenStudy (anonymous):

\[2x \sqrt{35x}\]

OpenStudy (anonymous):

How Did You do it

OpenStudy (anonymous):

Factor 140 first. Factor x^3 second.

OpenStudy (anonymous):

help please with that I know how 2 factor X^3

OpenStudy (anonymous):

140= 20 x 7 20= 4 x 5 4= 2 x 2 So you leave out the pair of twos. Then you multiply the remaining numbers and put them inside the radical sign. X^3= x^2 times x X^2= x times x So apply the same concept here. Leave out the pair of xs and put the third x under the radical sign